# SnapMoGen: Human Motion Generation from Expressive Texts

### v1

@@ Abstract
footnotetext: Project lead: cguo2@snapchat.com; bzhou@snapchat.com
@@ 1 Introduction
The importance of expressive text annotations has been well-established in other text-conditioned visual content synthesis fields betker2023improving ; manas2024improving ; hao2023optimizing ; chen2024tailored ; datta2023prompt . Descriptive prompts notably enhance the accuracy and aesthetic quality of generated images manas2024improving ; betker2023improving and improve temporal coherence in video generation bao2024vidu . These models understand complex visual content compositions by learning from rich text semantics, enabling fine-grained visual content editing kawar2023imagic ; chang2023muse , adaptation po2024orthogonal ; ruiz2023dreambooth ; sohn2023styledrop , and understanding chen2023videollm ; zhang2023video ; lin2023video . More importantly, when dealing with casual user prompts, rich LLM knowledge can be leveraged to enhance prompts with specific details and nuances that models have learned from fine-grained textual training data, effectively improving generalization capability. To foster this research direction in motion synthesis, we introduce SnapMoGen, which features high-quality motions captioned with accurate and highly expressive descriptions. SnapMoGen is created by segmenting long motion sequences into meaningful 4-12 second clips, each accompanied by six text descriptions—two manually annotated by human experts and four augmented by an LLM that introduces diversity while preserving semantics and temporal consistency. In total, SnapMoGen comprises 20K motion clips, amounting to 40 hours of mocap data, accompanied by 122K detailed text descriptions. As shown in Fig. 1, our text annotations contain extremely rich semantic cues of human movements, with an average length of 48 words—three times longer than HumanML3D. Furthermore, our continuous motion segments facilitate research in long-term motion synthesis and motion localization. Tab. 1 presents a statistical comparison between SnapMoGen and related motion-text datasets.
@@
To generate motions from expressive texts, we build an improved model upon the previous state-of-the-art approach—MoMask guo2024momask . MoMask applies residual quantization to motion latent features, transforming them into multiple ordered sets of same-length discrete token sequences. Despite achieving pleasing VQ reconstruction through extensive tokens, many of them are not utilized to their full capacity. For instance, tokens following the first quantization layer carry only marginal information. This inefficiency, combined with its layer-specific token vocabulary design, creates inflexibility in subsequent text-to-token generation—necessitating separate models for different token sequences: a primary model for the first sequence and a secondary model for remaining tokens. To overcome these limitations, we adopt a multi-scale approach for motion tokenization and generate all motion tokens using a single generative masked transformer. In our residual VQ, tokens at each quantization layer focus on a particular temporal scale, following a coarse-to-fine gradual progression. Additionally, we share one codebook across all layers to ensure a universal token vocabulary. As shown in Footnote 3, our multi-scale RVQ continually learns meaningful semantics with more layers, outperforming conventional RVQ guo2024momask with $45\%$ less tokens. Then, we simply concatenate all tokens along the temporal dimension and train a generative transformer to produce tokens from text by predicting randomly masked tokens. Our new framework, dubbed MoMask++, outperforms MoMask on text-to-motion generation with only a quarter of their token count, as in Tab. 3.
@@
Datasets Year # Clips Duration # Texts Avg. words per text Avg. length per clip Mocap? Continuous?
@@ 3 SnapMoGen Dataset
SnapMoGen encompasses 43.7 hours of high-quality motion data captured at 30 frames per second. The dataset comprises a total of 4.7M motion frames, featuring a diverse range of actions including daily activities, fitness routines, social interactions, dances, and more. We deliberately incorporate various stylized performances (e.g., princess, elderly person, zombie) to enhance diversity. SnapMoGen captures performances from 10 participants, resulting in 20,450 motion clips ranging from 4 to 12 seconds in length. Each motion clip is accompanied by 6 detailed textual descriptions (2 manually annotated, 4 LLM-augmented), totaling 122,565 textual descriptions with an average length of 48 words. A comparison between SnapMoGen and existing motion-text datasets is presented in Table 1. We further augment the dataset by mirroring motion data guo2022generating throughout our experiments.
@@ 4.1 Preliminary: Motion Tokenization via Residual Quantization
$\hat{f}=\texttt{lookup}(\mathcal{C},q),\quad\quad\quad\hat{\mathbf{m}}=\mathcal{D}(\hat{f}).$ (2)
@@
Each quantization layer $\mathcal{Q}^{v}(\cdot)$ contains a separate codebook $\mathcal{C}^{v}\in\mathbb{R}^{K\times d}$. This approach yields $V+1$ discrete token sequences $[q^{v}]_{0}^{V}\in[K]^{(V+1)\times n}$ of length $n$ for a motion sequence. The final approximation of the latent sequence $f$ is the sum of all quantized features $\hat{f}=\sum_{v=0}^{V}\hat{f}^{v}$.
@@ 4.2.1 Multi-scale Motion Quantization
As depicted in Fig. 2 (a), for a motion latent feature sequence $f\in\mathbb{R}^{n\times d}$, our quantizer employs a series of residual quantization operations at progressively increasing temporal resolutions $\{h^{v}\}_{v=0}^{V}$, where $h^{0}<\cdots<h^{V}=n$. We denote all quantization operations uniformly as $\mathcal{Q}(\cdot)$ as they share a common codebook $\mathcal{C}$. Quantized features at coarse level $\hat{f}^{v}\in\mathbb{R}^{h_{v}\times d}$ are bilinearly interpolated to full resolution $\hat{f}_{\uparrow}^{v}=\mathcal{I}(\hat{f}^{v},h^{V})$, where residuals are calculated and then downsampled to next scale ($h_{v+1}$) to be quantized by the succeeding layer. Mathematically, Eq. 3 is reformulated as:
@@
$\hat{f}=\sum_{v=0}^{V}\mathcal{I}(\hat{f}^{v},h^{V}),\quad\quad\,\,\,\hat{\mathbf{m}}=\mathcal{D}(\hat{f}).$
We further add self-attention layers after each res-block in existing motion VQVAE architectures guo2024momask ; zhang2023t2m for higher-fidelity motion reconstruction. During training, we introduce additional emphasis on reconstructing essential rotational features $\mathcal{L}_{ess}$ on the top of $\mathcal{L}_{rvq}$ in Eq. 4, weighed by $\lambda_{ess}$. The final learning objective becomes:
@@
As shown in Footnote 3, compared to previous all full-scale residual VQ guo2024momask , our multi-scale VQ effectively exploits token capacity and continually learns meaningful semantic features at each quantization layer. It achieves superior reconstruction quality significantly with fewer tokens (266 vs. 480).
@@ 4.2.2 Learning Text-to-token Mapping
We employ a single bidirectional transformer for token generation from text descriptions. Our framework is illustrated in Figure 2 (b-c). We utilize T5-base raffel2020exploring to extract word-level features from complex textual descriptions $c$. Motion tokens from all scales are concatenated along the temporal axis, yielding an extended token sequence $q$, which is then embedded through an MLP. We investigate two primary architectures for text conditioning: (i) In-context learning, where embeddings of motion tokens and text tokens are concatenated and processed uniformly as transformer input, and (ii) Cross-attention, which incorporates additional multi-head cross-attention layers that enable motion features to query relevant text features.
@@
MoMask++${}^{\text{in}}$ $\mathbf{{0.528}}^{\pm{.003}}$ $\mathbf{{0.718}}^{\pm{.003}}$ $\mathbf{{0.811}}^{\pm{.002}}$ ${0.072}^{\pm{.003}}$ $\mathbf{{2.912}}^{\pm{.008}}$ ${1.227}^{\pm{.046}}$
MoMask++${}^{\text{cra}}$ ${0.517}^{\pm{.002}}$ ${0.709}^{\pm{.002}}$ ${0.803}^{\pm{.002}}$ $\underline{{0.069}}^{\pm{.003}}$ ${2.948}^{\pm{.007}}$ ${1.192}^{\pm{.053}}$
@@
[…]
@@ 5 Experiments
We process motions in SnapMoGen following procedures established in HumanML3D, including motion mirroring and standardization. To prevent data leakage, we deliberately hold out a test (%10) set and a validation (%5) set where the motion scenarios (e.g., fashion) differ from the training motions. We primarily adopt the feature representation from HumanML3D, consisting of root angular velocity along Y-axis $\dot{r}^{a}\in\mathbb{R}$, root linear velocity on XZ-plane $\dot{r}^{xz}\in\mathbb{R}^{2}$, root height $\dot{r}^{y}\in\mathbb{R}$, 6D local joint rotations $\mathbf{j}^{r}\in\mathbb{R}^{6j}$, local joint positions $\mathbf{j}^{p}\in\mathbb{R}^{3j}$, and local joint velocities $\mathbf{j}^{v}\in\mathbb{R}^{3j}$, where $j$ denotes the number of joints. We empirically find that this comprehensive set of pose features leads to slightly better performance (Tab. 4). Our SnapMoGen follows a skeletal topology comprising 24 joints, resulting in 296-dimensional pose features. Unlike HumanML3D, our pose features are directly convertible to standard motion capture file formats (e.g., BVH).
@@
Our VQVAE encoder and decoder consists of three dilated res-blocks, with a down(up)-scale factor of 4 zhang2023t2m ; guo2024momask . The temporal quantization scales follows the progression $[n/2^{V},...,n/2^{0}]$ with $n$ denoting the full-scale length. We employ 4 (i.e., $V=3$) quantization layers for HumanML3D and 2 for SnapMoGen, with codebook sizes of $512\times 512$ and $2048\times 512$, respectively. The hyper-parameters $\beta$ and $\lambda_{ess}$ are set to 0.02 and 2.0. Our transformer architecture comprise 8 layers with feedforward size of 1024, latent dimension of 384, 6 attention heads, and a dropout ratio of 0.2, totaling 13.5M parameters for in-context model, and 18.3M parameters for cross-attention model. During inference, we use classifier-free guidance scales of 5 and 4, and iteration counts ($L$) of 10 and 18 for SnapMoGen and HumanML3D, respectively. All models are trained on a single Tesla V100 GPU, with batch size of 256 for VQVAEs and 64 for transformers.
@@
MoMask++${}^{\text{in}}$ $\mathbf{{0.805}}^{\pm{.002}}$ $\underline{{0.904}}^{\pm{.002}}$ $\mathbf{{0.938}}^{\pm{.001}}$ $\underline{{15.56}}^{\pm{.071}}$ $\underline{{0.684}}^{\pm{.001}}$ ${6.556}^{\pm{.178}}$
MoMask++${}^{\text{cra}}$ $\underline{{0.802}}^{\pm{.001}}$ $\mathbf{{0.905}}^{\pm{.002}}$ $\underline{{0.938}}^{\pm{.001}}$ $\mathbf{{15.06}}^{\pm{.065}}$ $\mathbf{{0.685}}^{\pm{.001}}$ ${7.259}^{\pm{.180}}$
@@ 5.1 Comparison to State-of-the-art Approaches
The quantitative results on HumanML3D and SnapMoGen  are reported in Tables 2 and 3, respectively. Overall, MoMask++ attains state-of-the-art performance on both datasets, demonstrating consistent improvements in motion-text alignment and motion quality. These advantages are particularly pronounced in our SnapMoGen dataset, partially due to the more expressive evaluation model. We observe that previous works struggle with the complex, lengthy text inputs in SnapMoGen, and fall short in maintaining multimodal semantic coherence, as evidenced by the relatively low CLIP scores and R-precision values. Notably, our method outperforms MoMask with only two VQ layers (a quarter of MoMask’s token count) with similar model size. Between the two variants of MoMask++, we find that the in-context model generally performs better on HumanML3D. It however tends to overfit on long text prompts in SnapMoGen (Fig. 5) and underperforms compared to the cross-attention model. Nevertheless, a significant gap to real motions still exists, suggesting substantial room for future improvements.
Figure 4 displays pose sequences generated by MoMask++, demonstrating its ability to produce precise motions following fine-grained text prompts. We further showcase the capability to handle out-of-domain user prompts by employing an LLM to rephrase the inputs. For additional generation results and comprehensive visual comparisons, please refer to the supplementary materials.
@@
Pose Dim. VQ Config. T2M Config. VQ Reconstruction T2M Generation
@@
base 296 2048 4 M In-context T5-base 384, 1024, 8 2.80 8.13 15.94 0.673
(A) 148 1024 4.71 7.12 15.95 0.667
(B) 1024 3.30 8.43 15.61 0.668
@@
3 3.13 8.40 16.21 0.652
5 2.31 7.65 16.38 0.663
(D) F 2.64 6.53 18.02 0.667
(E) 3.39 8.57 16.18 0.662
(F) CLIP 19.96 0.478
(G) 2 Cross-att. 384, 1024, 8 15.06 0.685
2 Cross-att. 512, 2048, 8 15.58 0.679
2 Cross-att. 512, 2048, 12 16.02 0.670
Table 4: Ablation analysis on SnapMoGen test set. "F/M" denotes full-scale versus multi-scale residual quantization. "Architecture" refers to transformer hyperparameters including latent dimension, feedforward size, and number of layers. Base: (384, 1024, 8). Medium: (512, 2048, 8). Large: (512, 2048, 12).

### v2

@@ Abstract
footnotetext: Project lead: guochuan5513@gmail.com; bzhou@snapchat.com
@@ 1 Introduction
The importance of expressive text annotations has been well-established in other text-conditioned visual content synthesis fields betker2023improving ; manas2024improving ; hao2023optimizing ; chen2024tailored ; datta2023prompt . Descriptive prompts notably enhance the accuracy and aesthetic quality of generated images manas2024improving ; betker2023improving and improve temporal coherence in video generation bao2024vidu . These models understand complex visual content compositions by learning from rich text semantics, enabling fine-grained visual content editing kawar2023imagic ; chang2023muse , adaptation po2024orthogonal ; ruiz2023dreambooth ; sohn2023styledrop , and understanding chen2023videollm ; zhang2023video ; lin2023video . More importantly, when dealing with casual user prompts, rich LLM knowledge can be leveraged to enhance prompts with specific details and nuances that models have learned from fine-grained textual training data, effectively improving generalization capability. To foster this research direction in motion synthesis, we introduce SnapMoGen, which features high-quality motions captioned with accurate and highly expressive descriptions. SnapMoGen is created by segmenting long motion sequences into meaningful 4-12 second clips, each accompanied by six text descriptions—two manually annotated by human experts and four augmented by an LLM that introduces diversity while preserving semantics and temporal consistency. In total, SnapMoGen comprises 20K motion clips, amounting to 40 hours of mocap data, accompanied by 122K detailed text descriptions. As shown in Fig.˜1, our text annotations contain extremely rich semantic cues of human movements, with an average length of 48 words—three times longer than HumanML3D. Furthermore, our continuous motion segments facilitate research in long-term motion synthesis and motion localization. Tab.˜1 presents a statistical comparison between SnapMoGen and related motion-text datasets.
@@
To generate motions from expressive texts, we build an improved model upon the previous state-of-the-art approach—MoMask guo2024momask . MoMask applies residual quantization to motion latent features, transforming them into multiple ordered sets of same-length discrete token sequences. Despite achieving pleasing VQ reconstruction through extensive tokens, many of them are not utilized to their full capacity. For instance, tokens following the first quantization layer carry only marginal information. This inefficiency, combined with its layer-specific token vocabulary design, creates inflexibility in subsequent text-to-token generation—necessitating separate models for different token sequences: a primary model for the first sequence and a secondary model for remaining tokens. To overcome these limitations, we adopt a multi-scale approach for motion tokenization and generate all motion tokens using a single generative masked transformer. In our residual VQ, tokens at each quantization layer focus on a particular temporal scale, following a coarse-to-fine gradual progression. Additionally, we share one codebook across all layers to ensure a universal token vocabulary. As shown in Footnote˜3, our multi-scale RVQ continually learns meaningful semantics with more layers, outperforming conventional RVQ guo2024momask with $45\%$ less tokens. Then, we simply concatenate all tokens along the temporal dimension and train a generative transformer to produce tokens from text by predicting randomly masked tokens. Our new framework, dubbed MoMask++, outperforms MoMask on text-to-motion generation with only a quarter of their token count, as in Tab.˜3.
@@
Datasets Year # Clips Duration # Texts
Avg. words
per text
Avg. length
per clip
Mocap? Continuous?
@@ 3 SnapMoGen Dataset
SnapMoGen encompasses 43.7 hours of high-quality motion data captured at 30 frames per second. The dataset comprises a total of 4.7M motion frames, featuring a diverse range of actions including daily activities, fitness routines, social interactions, dances, and more. We deliberately incorporate various stylized performances (e.g., princess, elderly person, zombie) to enhance diversity. SnapMoGen captures performances from 10 participants, resulting in 20,450 motion clips ranging from 4 to 12 seconds in length. Each motion clip is accompanied by 6 detailed textual descriptions (2 manually annotated, 4 LLM-augmented), totaling 122,565 textual descriptions with an average length of 48 words. A comparison between SnapMoGen and existing motion-text datasets is presented in Table˜1. We further augment the dataset by mirroring motion data guo2022generating throughout our experiments.
@@ 4.1 Preliminary: Motion Tokenization via Residual Quantization
[…]
@@
Each quantization layer $\mathcal{Q}^{v}(\cdot)$ contains a separate codebook $\mathcal{C}^{v}\in\mathbb{R}^{K\times d}$. This approach yields $V+1$ discrete token sequences $q^{v}]_{0}^{V}\in[K]^{(V+1)\times n}$ of length $n$ for a motion sequence. The final approximation of the latent sequence $f$ is the sum of all quantized features $\hat{f}=\sum_{v=0}^{V}\hat{f}^{v}$.
@@ 4.2.1 Multi-scale Motion Quantization
As depicted in Fig.˜2 (a), for a motion latent feature sequence $f\in\mathbb{R}^{n\times d}$, our quantizer employs a series of residual quantization operations at progressively increasing temporal resolutions $\{h^{v}\}_{v=0}^{V}$, where $h^{0}<\cdots<h^{V}=n$. We denote all quantization operations uniformly as $\mathcal{Q}(\cdot)$ as they share a common codebook $\mathcal{C}$. Quantized features at coarse level $\hat{f}^{v}\in\mathbb{R}^{h_{v}\times d}$ are bilinearly interpolated to full resolution $\hat{f}_{\uparrow}^{v}=\mathcal{I}(\hat{f}^{v},h^{V})$, where residuals are calculated and then downsampled to next scale ($h_{v+1}$) to be quantized by the succeeding layer. Mathematically, Eq.˜3 is reformulated as:
@@
[…]
We further add self-attention layers after each res-block in existing motion VQVAE architectures guo2024momask ; zhang2023t2m for higher-fidelity motion reconstruction. During training, we introduce additional emphasis on reconstructing essential rotational features $\mathcal{L}_{ess}$ on the top of $\mathcal{L}_{rvq}$ in Eq.˜4, weighed by $\lambda_{ess}$. The final learning objective becomes:
@@
As shown in Footnote˜3, compared to previous all full-scale residual VQ guo2024momask , our multi-scale VQ effectively exploits token capacity and continually learns meaningful semantic features at each quantization layer. It achieves superior reconstruction quality significantly with fewer tokens (266 vs. 480).
@@ 4.2.2 Learning Text-to-token Mapping
We employ a single bidirectional transformer for token generation from text descriptions. Our framework is illustrated in Figure˜2 (b-c). We utilize T5-base raffel2020exploring to extract word-level features from complex textual descriptions $c$. Motion tokens from all scales are concatenated along the temporal axis, yielding an extended token sequence $q$, which is then embedded through an MLP. We investigate two primary architectures for text conditioning: (i) In-context learning, where embeddings of motion tokens and text tokens are concatenated and processed uniformly as transformer input, and (ii) Cross-attention, which incorporates additional multi-head cross-attention layers that enable motion features to query relevant text features.
@@
[…]
@@
During inference, a complete sequence of $q$ can be generated in a constant number ($L$) of iterations. This process begins with an empty sequence $\left[\texttt{[MASK]}\right]^{N}$ where all tokens are masked, with $N=\sum_{v=0}^{V}h^{v}$ denoting the total number of tokens in $q$. At each iteration ($l$), the model predicts categorical token distributions at masked locations, samples tokens, and re-masks the $\lceil\gamma(\frac{l}{L}\cdot N)\rceil$ lowest-confidence tokens. This process repeats until $l$ reaches $L$. We also adopt classifier-free guidance as in  guo2024momask with guidance scale $s$. Finally, all generated tokens are decoded and projected back to motion sequence through the VQ-VAE decoder.
@@ 5 Experiments
We process motions in SnapMoGen following procedures established in HumanML3D, including motion mirroring and standardization. To prevent data leakage, we deliberately hold out a test (%10) set and a validation (%5) set where the motion scenarios (e.g., fashion) differ from the training motions. We primarily adopt the feature representation from HumanML3D, consisting of root angular velocity along Y-axis $\dot{r}^{a}\in\mathbb{R}$, root linear velocity on XZ-plane $\dot{r}^{xz}\in\mathbb{R}^{2}$, root height $\dot{r}^{y}\in\mathbb{R}$, 6D local joint rotations $\mathbf{j}^{r}\in\mathbb{R}^{6j}$, local joint positions $\mathbf{j}^{p}\in\mathbb{R}^{3j}$, and local joint velocities $\mathbf{j}^{v}\in\mathbb{R}^{3j}$, where $j$ denotes the number of joints. We empirically find that this comprehensive set of pose features leads to slightly better performance (Tab.˜4). Our SnapMoGen follows a skeletal topology comprising 24 joints, resulting in 296-dimensional pose features. Unlike HumanML3D, our pose features are directly convertible to standard motion capture file formats (e.g., BVH).
@@
Our VQVAE encoder and decoder consists of three dilated res-blocks, with a down(up)-scale factor of 4 zhang2023t2m ; guo2024momask . The temporal quantization scales follows the progression $n/2^{V},...,n/2^{0}]$ with $n$ denoting the full-scale length. We employ 4 (i.e., $V=3$) quantization layers for HumanML3D and 2 for SnapMoGen, with codebook sizes of $512\times 512$ and $2048\times 512$, respectively. The hyper-parameters $\beta$ and $\lambda_{ess}$ are set to 0.02 and 2.0. Our transformer architecture comprise 8 layers with feedforward size of 1024, latent dimension of 384, 6 attention heads, and a dropout ratio of 0.2, totaling 13.5M parameters for in-context model, and 18.3M parameters for cross-attention model. During inference, we use classifier-free guidance scales of 5 and 4, and iteration counts ($L$) of 10 and 18 for SnapMoGen and HumanML3D, respectively. All models are trained on a single Tesla V100 GPU, with batch size of 256 for VQVAEs and 64 for transformers.
@@
[…]
@@ 5.1 Comparison to State-of-the-art Approaches
The quantitative results on HumanML3D and SnapMoGen  are reported in Tables˜2 and 3, respectively. Overall, MoMask++ attains state-of-the-art performance on both datasets, demonstrating consistent improvements in motion-text alignment and motion quality. These advantages are particularly pronounced in our SnapMoGen dataset, partially due to the more expressive evaluation model. We observe that previous works struggle with the complex, lengthy text inputs in SnapMoGen, and fall short in maintaining multimodal semantic coherence, as evidenced by the relatively low CLIP scores and R-precision values. Notably, our method outperforms MoMask with only two VQ layers (a quarter of MoMask’s token count) with similar model size. Between the two variants of MoMask++, we find that the in-context model generally performs better on HumanML3D. It however tends to overfit on long text prompts in SnapMoGen (Fig.˜6) and underperforms compared to the cross-attention model. Nevertheless, a significant gap to real motions still exists, suggesting substantial room for future improvements.
Figure˜4 displays pose sequences generated by MoMask++, demonstrating its ability to produce precise motions following fine-grained text prompts. We further showcase the capability to handle out-of-domain user prompts by employing an LLM to rephrase the inputs. For additional generation results and comprehensive visual comparisons, please refer to the supplementary materials.
@@
Pose Dim. VQ Config. VQ Reconstruction T2M Generation
@@
base 296 2048 4 M 2.80 8.13 15.94 0.673
(A) Only essential pose feat. 148 1024 4.71 7.12 15.95 0.667
(B) Smaller codebook 1024 3.30 8.43 15.61 0.668
@@
1 8.81 10.48 16.25 0.677
(D) Full-scale v.s. multi-scale F 2.64 6.53 18.02 0.667
(E) W/o attention 3.39 8.57 16.18 0.662
Table 4: Ablation analysis of VQ configuration on SnapMoGen test set. "F/M" denotes full-scale versus multi-scale residual quantization. For text2motion, this experiment use 284 latent dimension, 1024 forward size, and 8 attention layers, with T5-base text encoder and in-context conditioning.
Text Aug. T2M Config. T2M Generation
[…]
base T5-base In-context (B) 384, 1024, 8 15.56 0.684
(A) W/o text aug. 17.98 0.656
(B) CLIP text enc. CLIP 19.96 0.478
(C) Cross-att cond. Cross-att. 15.06 0.685
(D) Larger model Cross-att. (M) 512, 2048, 8 15.58 0.679
Cross-att. (L) 512, 2048, 12 16.02 0.670
Table 5: Ablation analysis of T2M model configuration on SnapMoGen test set. "Architecture" refers to transformer hyperparameters including latent dimension, feedforward size, and number of layers. This experiment use 2 quantization layers for VQ, with 2048 codebook size and 296 pose features.